OpenStudy (anonymous):

Evaluate the expression. log_{2}72-log_{2}9

OpenStudy (anonymous):

\[\log_{2}72-\log_{2}9 \]

OpenStudy (anonymous):

3

OpenStudy (anonymous):

log2 (72/9) = log2 (8) = log 2 (2^3) = 3